# Wasserstein Spatial Depth

### v1

@@ 8.1 Consistency of the empirical Wasserstein spatial depth
The simulation results below support the theoretical results in Section 6. That is, the empirical WSD, formulated in (9), is close to the theoretical value $\mathrm{SD}(Q;{\bf P})$ in Definition 3.1. Hence, the WSD can be inferred accurately from sample data and has practical value. Four cases are considered and described below.
@@
  • Case 2: ${\bf P}$ is supported on a family of Weibull distributions with fixed scale parameter $1$ and varying shape parameter $k$. This family of Weibull distributions is indexed by the shape parameter $k$ which takes value either $1$ or $2$ with equal probabilities, i.e., $k\sim\mathrm{Unif}\{1,2\}$. Let $Q$ be the Weibull distribution with shape parameter $k_{Q}$ ($k_{Q}$ equating either $1$ or $2$). Its theoretical WSD with respect to ${\bf P}$ is
    $\mathrm{SD}(Q;{\bf P})$ $=1-\sqrt{\int_{0}^{\infty}\Big{(}\mathbb{E}_{k\sim\mathrm{Unif}\{1,2\}}\frac{x-x^{k_{Q}/k}}{\mathcal{W}_{2}(k_{Q},k)}\Big{)}^{2}k_{Q}x^{k_{Q}-1}e^{-x^{k_{Q}}}\mathrm{d}x}$
@@
    using $\overline{k}_{Q}=3-k_{Q}$, the convention $/0=0$ 0, and where $\mathcal{W}_{2}(k_{Q},k)$ is derived by
    $\mathcal{W}_{2}(k_{Q},k)$ $=\sqrt{\int_{0}^{\infty}\big{(}x-x^{k_{Q}/k}\big{)}^{2}k_{Q}x^{k_{Q}-1}e^{-x^{k_{Q}}}\mathrm{d}x}.$
  • Case 3: ${\bf P}$ is supported on a family of isotropic bivariate Gaussian distributions with varying centers. The distribution of the Gaussian centers is supported on four points $\{\boldsymbol{\mu}_{1}=(1,0)^{\top},\boldsymbol{\mu}_{2}=(-1,0)^{\top},\boldsymbol{\mu}_{3}=(0,1)^{\top},\boldsymbol{\mu}_{4}=(0,-1)^{\top}\}$ with equal probabilities $1/4$. Let $Q$ be any one of $\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{I}_{2\times 2})$ for $k=1,\ldots,4$. The theoretical WSD is computed as, see Section 4.3,
@@
  • Case 4: ${\bf P}$ is supported on a family of bivariate uniform distributions $\mathrm{Unif}\big{(}[0,c]^{2}\big{)}$ with $c\sim\mathrm{Unif}[1,2]$. Let $Q$ be $\mathrm{Unif}\big{(}[0,c_{q}]^{2}\big{)}$. Its theoretical WSD with respect to ${\bf P}$ is
    $\mathrm{SD}(Q;{\bf P})$ $=1-\sqrt{\int_{[0,c_{q}]^{2}}\bigg{\|}\mathbb{E}_{c\sim\mathrm{Unif}[1,2]}\ \frac{\boldsymbol{x}-(c/c_{q})\boldsymbol{x}}{\mathcal{W}_{2}(c_{q},c)}\bigg{\|}^{2}\frac{1}{c_{q}^{2}}\mathrm{d}\boldsymbol{x}}$
    $=1-\sqrt{\int_{[0,c_{q}]^{2}}\bigg{\|}\sqrt{3/2}\boldsymbol{x}\mathbb{E}_{c\sim\mathrm{Unif}[1,2]}\frac{1-(c/c_{q})}{|c_{q}-c|}\bigg{\|}^{2}\frac{1}{c_{q}^{2}}\mathrm{d}\boldsymbol{x}}$
@@
Figure 1: The green solid line depicts the change of theoretical WSD along the parameter indexing ${\bf P}$. The black circles represent the distribution of empirical WSDs, with error bars indicating one standard deviation above and below the mean
@@
  • Case 1: $\mathbf{P}$ is supported on a set of $d=10$ dimensional Gaussian distributions with identity covariance matrix. The Gaussian centers are i.i.d. from $\mathrm{Unif}[-2,2]^{d}$.
  • Case 2: $\mathbf{P}$ is supported on a set of $d=10$ dimensional Gaussian distributions with a common covariance matrix. The common covariance matrix is chosen as $\Sigma_{i,j}=0.2^{|i-j|}$. The Gaussian centers are drawn in the same way as Case 1.
  • Case 3: The support of $\mathbf{P}$ is a set of uniform distributions on $d=10$ dimensional unite cubes with varying centers. The centers of the cubes are identically independently drawn from $\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{d\times d})$.
@@
The simulation procedure is as follows. First, $n=500$ distributions are drawn as described above in each case. Second, $m=500$ data points are randomly drawn for each sampled distribution. Third, the empirical WSD of each empirical distribution is computed according to (9), and the empirical spatial depths of the locations are computed as in (5). Finally, we check whether the empirical WSDs and spatial depths are approximately equal. As shown in Figure 2, there are nice equality relationships between the two depths.
Figure 2: The relationships between the WSD and conventional spatial depth in the four cases above.
@@ 8.3 Outlier detection
    $\big{[}\mathrm{Gamma}(3,2)\big{]}^{d},\quad\big{[}\mathrm{Unif}(-6,6)\big{]}^{d},\quad\big{[}\mathrm{Beta}(0.1,0.1)\big{]}^{d},$
@@
  • Case 2: the population is a collection of uniform distributions $\big{[}\mathrm{Unif}(0,u)\big{]}^{d}$ with $u\sim\mathrm{Unif}(1,2)$; the outlier distributions are
@@
For each case, we repeat similar experiments for $20$ times. The experimental procedure is as follows: draw the data array $\boldsymbol{X}$ according to Case 1 or Case 2; compute their empirical WSD according to (9); detect the outlier distributions whose empirical WSDs are smaller than the $1\%$ quantile of all the empirical WSDs. In each of these $20$ repetitions, all the outlier distributions can be detected for both Case 1 and Case 2. Figure 3 shows the result of one experiment, where the black and red dots represent distributions from ${\bf P}$ and the outlier distributions, respectively.
@@
Figure 3: Left panel: the distributions are drawn according to Case 1. Right panel: the distributions are drawn according to Case 2. The red dots represent the outlier distributions.
@@ 9 Application
We consider monthly temperatures obtained by averaging daily temperatures per month. To each weather station, we obtain a $12$ monthly-average temperature curve, represented by a vector in $\mathbb{R}^{12}$. Hence, the monthly temperatures of each year correspond to one distribution on $\mathbb{R}^{12}$. For a particular year, the $12$ monthly temperatures (forming a vector in $\mathbb{R}^{12}$) collected at each station act as a sample point drawn from this distribution. Finally, we gather $150$ distributions (from year 1874 to year 2023) with each distribution associated to $40$ sample points (for the $40$ meteorological stations), and where each sample point is a $12$-dimensional real vector. Contrary to other work, we do not consider the annual evolution of the temperatures for a particular place but rather analyze the different temperature curves at all locations at the same time. We aim at understanding weather change at a global scale by considering the $40$ different locations as representatives of the European climate.
Within this framework, we compute the empirical WSDs of these $150$ distributions as in (9). Several outlier years are identified based on their excessively small WSDs. As we discuss next, these identified “abnormal years” are consistent with historical records, which further validate the practical utility of Wasserstein spatial depth.
@@
Figure 4: The green dots represent regular/representative/central distributions, while the red dots correspond to the distributions near the outskirt and “far” from the center
The values of the $150$ empirical WSDs are shown in Figure 4. The lowest $5\%$ values, which we consider as outliers, are colored red, and the corresponding years are also marked. Based on empirical WSDs, the temperatures at years 1879, 1929, 1940, 1942, 1947, 1956, 1963, and 2018 are more “exotic” or near outskirt. After searching among historical documentations, we indeed fond evidences to support this discovery. Year 1879 was an extremely cold year, featured with a unusually snowy winter (November and December); the first two months of 1929 were recorded as one of the coldest winters in Europe during the past century with temperature reaching down to -30°C in central Europe. Both year 1940 and year 1942 were marked by severe winters with dramatic ice storms, and year 1942 had a cool summer; The weather in year 1947 was unusually cold in winter and record-breaking hot in summer; Europe experienced severe cold waves in both winters of 1956 and 1963; The well-known 2018 European drought and heat wave led to record-breaking temperatures and wildfires in many parts of Europe.
@@
Figure 5: Comparisons between the most regular years 1935 and 1960 (the two years with the largest Wasserstein spatial depths) and four outlier years. In each plot, the bundle of green curves represents the temperature trends in years 1935 and 1960 at 40 locations in Europe (totally $80$ green curves); the bundle of red curves represents the temperature trends in an outlier year at the same 40 locations (totally $40$ red curves).
@@ 10 Further directions and future work
In this work, we propose a new notion of depth on the Wasserstein space. We demonstrate that it preserves critical properties of conventional statistical depths. Additionally, it has a straightforward empirical counterpart that can be easily computed from sample data and is asymptotically consistent. Numerical simulations and real data analysis further support its practical utility.
Note that we have defined the new notion of Wasserstein spatial depth $\mathrm{SD}(Q;{\bf P})$ for absolutely continuous distributions $Q$ and where ${\bf P}$ can be arbitrary. This is because our approach exploits the definition of the geodesics in the Wasserstein space (see Section 3.3).
When $Q$ is not absolutely continuous, the geodesic between $Q$ and another distribution $P$ might be not unique. In this case, the set of geodesics is given by the laws of the random vectors $(1-t)\mathbf{X}+t\mathbf{Y}$ where the law of the random vector $(\mathbf{X},\mathbf{Y})$, namely $\pi_{P,Q}$, is an optimal transport plan, as in (1) with $p=2$. Hence, uniqueness of the geodesics is equivalent to uniqueness of the transport plans.
@@
We leave for future exploration the practical utility of this complementary WSD, along with the task of establishing analogous favorable mathematical properties as those demonstrated in this paper. Note that the depth in (10) coincides with $\mathrm{SD}(Q;{\bf P})$ in the special case where both $Q$ and (a.s.) the samples from P are absolutely continuous. This can be seen from the arguments leading to (A.3).
Finally, for computational reasons, the statistics and machine learning community has also focused on regularized optimal transport [15, 35]. It is an interesting prospect as well to extend the Wasserstein spatial depth to regularized optimal transport.

### v2

@@ 8.1 Consistency of the empirical Wasserstein spatial depth
The simulation results below support the theoretical results in Section 6. That is, the empirical WSD, formulated in (8.1), is close to the theoretical value $\mathrm{SD}(Q;{\bf P})$ in Definition 3.1. Hence, the WSD can be inferred accurately from sample data and has practical value. Four cases are considered and described below.
@@
  • Case 2: ${\bf P}$ is supported on a family of Weibull distributions with fixed scale parameter $1$ and varying shape parameter $k$. This family of Weibull distributions is indexed by the shape parameter $k$ which takes value either $1$ or $2$ with equal probabilities, i.e., $k\sim\mathrm{Unif}(\{1,2\})$. Let $Q$ be the Weibull distribution with shape parameter $k_{Q}$ ($k_{Q}$ equating either $1$ or $2$). Its theoretical WSD with respect to ${\bf P}$ is
    $\mathrm{SD}(Q;{\bf P})$ $=1-\sqrt{\int_{0}^{\infty}\Big{(}\mathbb{E}_{k\sim\mathrm{Unif}(\{1,2\})}\frac{x-x^{k_{Q}/k}}{\mathcal{W}_{2}(k_{Q},k)}\Big{)}^{2}k_{Q}x^{k_{Q}-1}e^{-x^{k_{Q}}}\mathrm{d}x}$
@@
    using $\overline{k}_{Q}=3-k_{Q}$, the convention $/0=0$ 0, and where $\mathcal{W}_{2}(k_{Q},\overline{k}_{Q})$ is derived by
    $\mathcal{W}_{2}(k_{Q},\overline{k}_{Q})$ $=\sqrt{\int_{0}^{\infty}\big{(}x-x^{k_{Q}/\overline{k}_{Q}}\big{)}^{2}k_{Q}x^{k_{Q}-1}e^{-x^{k_{Q}}}\mathrm{d}x}.$
  • Case 3: ${\bf P}$ is supported on a family of isotropic bivariate Gaussian distributions with varying centers. The distribution of the Gaussian centers is supported on four points $\{\boldsymbol{\mu}_{1}=(1,0)^{\top},\boldsymbol{\mu}_{2}=(-1,0)^{\top},\boldsymbol{\mu}_{3}=(0,1)^{\top},\boldsymbol{\mu}_{4}=(0,-1)^{\top}\}$ with equal probabilities $1/4$. Let $Q$ be $\mathcal{N}(\boldsymbol{\mu}_{q},\boldsymbol{I})$ for $q\in\{1,\ldots,4\}$. The theoretical WSD is computed as, see Section 4.3,
@@
  • Case 4: ${\bf P}$ is supported on a family of bivariate uniform distributions $\mathrm{Unif}\big{(}[0,c]^{2}\big{)}$ with $c\sim\mathrm{Unif}([1,2])$. Let $Q$ be $\mathrm{Unif}\big{(}[0,c_{q}]^{2}\big{)}$. Its theoretical WSD with respect to ${\bf P}$ is
    $\mathrm{SD}(Q;{\bf P})$ $=1-\sqrt{\int_{[0,c_{q}]^{2}}\bigg{\|}\mathbb{E}_{c\sim\mathrm{Unif}([1,2])}\ \frac{\boldsymbol{x}-(c/c_{q})\boldsymbol{x}}{\mathcal{W}_{2}(c_{q},c)}\bigg{\|}^{2}\frac{1}{c_{q}^{2}}\mathrm{d}\boldsymbol{x}}$
    $=1-\sqrt{\int_{[0,c_{q}]^{2}}\bigg{\|}\sqrt{3/2}\boldsymbol{x}\mathbb{E}_{c\sim\mathrm{Unif}([1,2])}\frac{1-(c/c_{q})}{|c_{q}-c|}\bigg{\|}^{2}\frac{1}{c_{q}^{2}}\mathrm{d}\boldsymbol{x}}$
@@
Figure 1: The green solid lines depict the change of theoretical WSD along the parameter indexing ${\bf P}$. The black circles represent the distribution of empirical WSDs, with error bars indicating one standard deviation above and below the mean.
@@
  • Case 1: $\mathbf{P}$ is supported on a set of $d=10$-dimensional Gaussian distributions with identity covariance matrix. The Gaussian centers are i.i.d. from $\mathrm{Unif}([-2,2]^{d})$.
  • Case 2: $\mathbf{P}$ is supported on a set of $d=10$-dimensional Gaussian distributions with a common covariance matrix. The common covariance matrix is chosen as $\Sigma_{i,j}=0.2^{|i-j|}$. The Gaussian centers are drawn in the same way as in Case 1.
  • Case 3: The support of $\mathbf{P}$ is a set of uniform distributions on $d=10$-dimensional unit cubes with varying centers. The centers of the cubes are identically independently drawn from $\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$.
@@
The simulation procedure is as follows. First, $n=500$ distributions are drawn as described above in each case. Second, $m=500$ data points are randomly drawn for each sampled distribution. Third, the empirical WSD of each empirical distribution is computed according to (8.1), and the empirical spatial depths of the locations are computed as in (4.1). Finally, we check whether the empirical WSDs and spatial depths are approximately equal. As shown in Figure 2, there are nice equality relationships between the two depths.
Figure 2: The relationships between the WSD and conventional spatial depth in the four cases of Section 8.2.
@@ 8.3 Outlier detection
    $\big{[}\mathrm{Gamma}(3,2)\big{]}^{d},\quad\big{[}\mathrm{Unif}([-6,6])\big{]}^{d},\quad\big{[}\mathrm{Beta}(0.1,0.1)\big{]}^{d},$
@@
  • Case 2: the population is a collection of uniform distributions $\big{[}\mathrm{Unif}([0,u])\big{]}^{d}$ with $u\sim\mathrm{Unif}([1,2])$; the outlier distributions are
@@
For each case, we repeat similar experiments for $20$ times. The experimental procedure is as follows: draw the data array $\boldsymbol{X}$ according to Case 1 or Case 2; compute their empirical WSD according to (8.1); detect the outlier distributions whose empirical WSDs are smaller than the $1\%$ quantile of all the empirical WSDs. In each of these $20$ repetitions, all the outlier distributions can be detected for both Case 1 and Case 2. Figure 3 shows the result of one experiment, where the black and orange dots represent distributions from ${\bf P}$ and the outlier distributions, respectively.
@@
Figure 3: Left panel: the distributions are drawn according to Case 1. Right panel: the distributions are drawn according to Case 2. The green dots represent regular distributions from the population $\mathbf{P}$, and the orange dots represent the outlier distributions.
8.4 Wasserstein spatial depth vs. functional depth
Hilbertian embedding of probability measures (into a RKHS) is a powerful technique in machine learning and statistics that allows for a functional representation of probability measures [52]. This approach maps a probability distribution $\mu\in\mathcal{P}(\mathbb{R}^{d})$ to an element $f_{\mu}$ in the RKHS $\mathcal{H}_{K}$ via kernel mean embedding,
[…]
Here $K:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ is a kernel on $\mathbb{R}^{d}$, yielding the Hilbert space $\mathcal{H}_{K}$ (the RKHS) of functions from $\mathbb{R}^{d}$ to $\mathbb{R}$ [4]. Given this embedding machinery and an available notion of depth for functional data [27, 38], one could first transform a distribution into a functional data point and then compute its functional depth. However, such an approach neglects the rich geodesic structure of the Wasserstein space. The relative distance and “ordering” of pairs of distributions are probably distorted after Hilbertian embedding. The simulation results in this subsection support the above point of view.
We consider two cases here. In each case, $n=100$ similar distributions (denoted as regular distributions) and four exotic distributions are drawn. By “similar” we mean that these $n$ distributions are of the same parametric family and are close to each other in terms of Wasserstein distance. All the distributions are on $\mathbb{R}^{3}$ so that visualization is possible. We draw $m=300$ data points for each distribution.
  • Case 1: the regular distributions are from a collection of spherical Gaussian distributions $N(\boldsymbol{\mu},\sigma^{2}\boldsymbol{I})$ with varying centers $\boldsymbol{\mu}\stackrel{{\scriptstyle\textit{i.i.d.}}}{{\sim}}\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ and varying variances $\sigma\stackrel{{\scriptstyle\textit{i.i.d.}}}{{\sim}}\mathrm{Unif}([0.8,1])$; the four exotic distributions are
    $\big{[}\mathrm{Gamma}(3,2)\big{]}^{d},\quad\big{[}\mathrm{Weibull}(2,1)\cdot 3\mathrm{Bernoulli}(-1,1,1/2)\big{]}^{d},$
    $\big{[}\mathrm{Unif}(\{-3.5,-2.5,2.5,3.5\})\big{]}^{d},\quad\mathcal{N}((-3,3,-3)^{\top},\boldsymbol{\Sigma})\ \text{with}\ \Sigma_{i,j}=0.5^{|i-j|}.$
  • Case 2: the regular distributions are from a collection of uniform distributions $\big{[}\mathrm{Unif}([0,u])\big{]}^{d}$ with $u\sim\mathrm{Beta}(2,2)+1$; the exotic distributions are
    $\big{[}\mathrm{Poisson}(1)\big{]}^{d},\quad\big{[}\mathrm{Exponential}(2)\cdot\mathrm{Bernoulli}(-1,1,1/2)\big{]}^{d},$
    $\big{[}\mathrm{Unif}(\{1,2,3\})\big{]}^{d},\quad\mathrm{Multinomial}\big{(}2d,(0.1,0.2,0.7)\big{)}.$
Here $\mathrm{Bernoulli}(-1,1,1/2)$ means an independent Bernoulli random variable taking value $-1$ or $+1$ with probability $1/2$. In each case, the regular distributions are close to each other in the Wasserstein space because
$\mathcal{W}_{2}\Big{(}N(\boldsymbol{\mu}_{1},\sigma_{1}^{2}\boldsymbol{I}),N(\boldsymbol{\mu}_{2},\sigma_{2}^{2}\boldsymbol{I})\Big{)}=\sqrt{\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|^{2}+d(\sigma_{1}-\sigma_{2})^{2}}\lesssim 1.5\sqrt{d},$
$\mathcal{W}_{2}\Big{(}\big{[}\mathrm{Unif}[0,u_{1}]\big{]}^{d},\big{[}\mathrm{Unif}[0,u_{2}]\big{]}^{d}\Big{)}$ $=\sqrt{\int_{[0,u_{1}]^{d}}\frac{\|\mathbf{x}-(u_{2}/u_{1})\mathbf{x}\|^{2}}{u_{1}^{d}}\mathrm{d}\mathbf{x}}$
$=\sqrt{d/3}\ |u_{2}-u_{1}|\leq\sqrt{d/3}.$
Also shown in Figure 4 (a) and Figure 5 (a), the regular distributions (represented by green triangles) tend to form a data cloud and are not visually distinguishable, while the exotic distributions are visually distant to the regular distributions.
We compare the WSD and two types of functional depth, Modified Band Depth (MBD) [38] and Functional Spatial Depth (FSD) [11] in terms of detecting those exotic distributions. To compute the functional depth of a distribution, we first embed the distribution into a RKHS via a Gaussian kernel, and then compute the functional depth of the embedded function. The MBD and FSD are computed, respectively, by the R packages depthTools and fda.usc. As shown in Figures 4 and 5, the WSD is able to detect the exotic distributions in both cases, while the functional depths are not informative on the “ordering” of the distributions. This numerical experiment shows the superiority of the proposed WSD when applied to distribution-valued data objects, which is expected since the WSD is specially designed for distribution-valued data objects and adapts to the geometry of the Wasserstein space.
Figure 4: (a): The data points are drawn from the distributions of Case 1. The green triangles represent data points from the regular distributions, while orange triangles represent data points from the exotic distributions. (b): The green dots represent the WSD values of the regular distributions, while the orange dots represent the WSD values of the exotic distributions. (c): Each dot represents the MBD of a distribution. The coloring pattern is the same as before. (d): Each dot represents the FSD of a distribution. The coloring pattern remains the same.
Figure 5: (a): The data points are drawn from the distributions of Case 2. The green triangles represent data points from the regular distributions, while orange triangles represent data points from the exotic distributions. (b): The green dots represent the WSD values of the regular distributions, while the orange dots represent the WSD values of the exotic distributions. (c): Each dot represents the MBD of a distribution. The coloring pattern is the same as before. (d): Each dot represents the FSD of a distribution. The coloring pattern remains the same.
@@ 9 Application
We consider monthly temperatures obtained by averaging daily temperatures per month. For each weather station, we obtain a $12$ monthly-average temperature curve, represented by a vector in $\mathbb{R}^{12}$. Hence, the monthly temperatures of each year correspond to one distribution on $\mathbb{R}^{12}$. For a particular year, the $12$ monthly temperatures (forming a vector in $\mathbb{R}^{12}$) collected at each station act as a sample point drawn from this distribution. Finally, we gather $150$ distributions (from year 1874 to year 2023) with each distribution associated to $40$ sample points (for the $40$ meteorological stations), and where each sample point is a $12$-dimensional real vector. In the following we assume that the distributions are drawn each year independently. Contrary to other work, we do not consider the annual evolution of the temperatures for a particular place but rather analyze the different temperature curves at all locations at the same time. We aim at understanding weather change at a global scale by considering the $40$ different locations as representatives of the European climate.
Within this framework, we compute the empirical WSDs of these $150$ distributions as in (8.1). Several outlier years are identified based on their excessively small WSDs. As we discuss next, these identified “abnormal years” are consistent with historical records, which further validate the practical utility of the WSD.
@@
Figure 6: The green dots represent regular/representative/central distributions, while the red dots correspond to the distributions near the outskirt and “far” from the center.
The values of the $150$ empirical WSDs are shown in Figure 6. The lowest $5\%$ values, which we consider as outliers, are colored red, and the corresponding years are also marked. Based on empirical WSDs, the temperatures at years 1879, 1929, 1940, 1942, 1947, 1956, 1963, and 2018 are more “exotic” or near outskirt. After searching among historical documentations, we indeed fond evidences to support this discovery. Year 1879 was an extremely cold year, featured with a unusually snowy winter (November and December). The first two months of 1929 were recorded as one of the coldest winters in Europe during the past century with temperature reaching down to -30°C in central Europe. Both year 1940 and year 1942 were marked by severe winters with dramatic ice storms, and year 1942 had a cool summer. The weather in year 1947 was unusually cold in winter and record-breaking hot in summer. Europe experienced severe cold waves in both winters of 1956 and 1963. The well-known 2018 European drought and heat wave led to record-breaking temperatures and wildfires in many parts of Europe.
@@
Figure 7: Comparisons between the most regular years 1935 and 1960 (the two years with the largest WSDs) and four outlier years. In each plot, the bundle of green curves represents the temperature trends in years 1935 and 1960 at 40 locations in Europe (totally $80$ green curves); the bundle of red curves represents the temperature trends in an outlier year at the same 40 locations (totally $40$ red curves).
@@ 10 Further directions and future work
In this work, we propose a new notion of depth on the Wasserstein space. We demonstrate that it preserves critical properties of conventional statistical depths. Additionally, it has a straightforward empirical counterpart that can be easily computed from sample data and is asymptotically consistent. Numerical simulations and real data analysis further support its practical utility. Importantly, in Section 8.4, we demonstrate that simply embedding distributions into linear Hilbert spaces, and relying on existing FDA methods, is not satisfying. In contrast, the WSD proves to be very informative in this section.
Note that we have defined the new notion of WSD $\mathrm{SD}(Q;{\bf P})$ for absolutely continuous distributions $Q$ and where ${\bf P}$ can be arbitrary. This is because our approach exploits the definition of the geodesics in the Wasserstein space (see Section 3.3).
When $Q$ is not absolutely continuous, the geodesic between $Q$ and another distribution $P$ might be not unique. In this case, the set of geodesics is given by the laws of the random vectors $(1-t)\mathbf{X}+t\mathbf{Y}$ where the law of the random vector $(\mathbf{X},\mathbf{Y})$, namely $\pi_{P,Q}$, is an optimal transport plan, as in (3.1) with $p=2$. Hence, uniqueness of the geodesics is equivalent to uniqueness of the transport plans.
@@
We leave for future exploration the practical utility of this complementary WSD, along with the task of establishing analogous favorable mathematical properties as those demonstrated in this paper. Note that the depth in (10.1) coincides with $\mathrm{SD}(Q;{\bf P})$ in the special case where both $Q$ and (a.s.) the samples from P are absolutely continuous. This can be seen from the arguments leading to (A.3) in the Appendix.
Finally, for computational reasons, the statistics and machine learning community has also focused on regularized optimal transport [21, 49]. It is an interesting prospect as well to extend the WSD to regularized optimal transport.